OpenStudy (anonymous):

A baseball team plays in a stadium that holds 51,000 spectators. With ticket prices at $10, the average attendance had been 38,000. When ticket prices were lowered to $8, the average attendance rose to 42,000. Find the demand function (price p as a function of attendance x), assuming it to be linear??

OpenStudy (anonymous):

I just need to find the equation. i keep getting this and its wrong : p = -1000 x+ 59,000

jimthompson5910 (jim_thompson5910):

With ticket prices at $10, the average attendance had been 38,000 means x = 38000 p = 10 giving you one point (38000, 10)

jimthompson5910 (jim_thompson5910):

When ticket prices were lowered to $8, the average attendance rose to 42,000 means x = 42000 p = 8 so another point is (42000, 8)

jimthompson5910 (jim_thompson5910):

now find the slope of the line through those two points

OpenStudy (anonymous):

yes and i get p = -1000 x+ 59,000

OpenStudy (anonymous):

it says its wrong

jimthompson5910 (jim_thompson5910):

the slope is way to big

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 8- 10)/(42000-38000) m = -2/4000 m = -1/2000 So the slope of the line that passes through the two points is -1/2000

OpenStudy (anonymous):

and how u find b?

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-1/2000)x + b 38000 = (-1/2000)(10) + b solve for b

OpenStudy (anonymous):

im getting b=7600001/200 n i believe thats way too big

jimthompson5910 (jim_thompson5910):

38000 = (-1/2000)(10) + b 38000 = -1/200 + b 38000 + 1/200 = b 7600000/200 + 1/200 = b 7600001/200 = b b = 7600001/200 b = 38000.005

OpenStudy (anonymous):

so y=(-1/2000)x +760001/200?

OpenStudy (anonymous):

it says its wrong :(

jimthompson5910 (jim_thompson5910):

try using decimal form instead of fraction form

OpenStudy (anonymous):

(a) Find the demand function (price p as a function of attendance x), assuming it to be linear.

OpenStudy (anonymous):

thats the actual question

jimthompson5910 (jim_thompson5910):

i know

jimthompson5910 (jim_thompson5910):

maybe you entered the fraction in wrong, try it in decimal form

OpenStudy (anonymous):

yea it doesn't let me put in decimals

jimthompson5910 (jim_thompson5910):

ok post a screenshot of what you're inputting

OpenStudy (anonymous):

sec

jimthompson5910 (jim_thompson5910):

k

jimthompson5910 (jim_thompson5910):

the numbers are different from what you gave me

OpenStudy (anonymous):

yea I just noticed wow.. im sorry

jimthompson5910 (jim_thompson5910):

it's fine, use the same idea just with these correct values

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

yw